OpenStudy (anonymous):

Prove that 38^3+37^3 is divisible by 75

OpenStudy (anonymous):

my first guess would be to factor

OpenStudy (anonymous):

\[a^3+b^3=(a+b)(a^2-ab+b^2)\] try it with \(a=38,b=37\)

OpenStudy (anonymous):

(38+37)(38^2-38*37+37^2)

geerky42 (geerky42):

And (38+37) = 75, thus it is divisible by 75 Q.E.D.